OpenStudy (anonymous):

solve simultaneous algebraically x-y=6 2x+3y=7

OpenStudy (anonymous):

time 3 to the top which equals, 3x-3y=18 then add both equations which then get one equation: 5x=25

OpenStudy (anonymous):

(3x-3y=18) +(2x+3y=7) =5x+0= 25 x=5

OpenStudy (anonymous):

or i guess you could solve for x or y in the top equation x=6+y then plug into the second equation

OpenStudy (anonymous):

x-y=6 2x+3y=7 x=6+y 2(6+y)+3y=7 12+5y=7 y=-5 x=1

OpenStudy (anonymous):

@Nick, you have a wrong calculation: \[12+5y=7 \Rightarrow 5y = -5 \Rightarrow y = -1\] then: \[x = 6+y \Rightarrow x = 6+(-1) \Rightarrow x = 5\]